\int \left(\frac{sin\left(x+a\right)}{m}+\frac{cos\left(x+b\right)}{n}\right)dx
1
Expert's answer
2020-10-05T17:07:40-0400

\int \left(\frac{sin\left(x+a\right)}{m}+\frac{cos\left(x+b\right)}{n}\right)dx"\int \left(\frac{sin\left(x+a\right)}{m}+\frac{cos\left(x+b\right)}{n}\right)dx=\int \left(\frac{sin\left(x+a\right)}{m}\right)dx+\int \left(\frac{cos\left(x+b\right)}{n}\right)dx="

"-\frac{cos\left(x+a\right)}{m}+\frac{sin\left(x+b\right)}{n}+C"

Answer: "\frac{sin\left(x+b\right)}{n}-\frac{cos\left(x+a\right)}{m}+C"